Marshalls Corporation has a bond currently outstanding. The bond has a face value of $1,000 and matures in 10 years. The bond makes no coupon payments for the first three years, then pays $25 every six months over the subsequent four years, and finally pays $85 every six months over the last three years. If the required return on these bonds is 5.5 percent compounded semiannually, what is the current price of the bond?

$986.33

$997.88

$1,049.55

$1,027.68

$1,009.82
Answer #1
Price of the bond is itspresent value of future cash flow
here payment is made semiannually hence we we should required return by 2
required reurn = 5.5% /2 =2.75%
Calculating present value
Year Cash Flow PV Factor PV Of Cash Flow
a b c=1/1.0275^a d=b*c
7 $                 25 0.827041 $                   20.68
8 $                 25 0.804906 $                   20.12
9 $                 25 0.783364 $                   19.58
10 $                 25 0.762398 $                   19.06
11 $                 25 0.741993 $                   18.55
12 $                 25 0.722134 $                   18.05
13 $                 25 0.702807 $                   17.57
14 $                 25 0.683997 $                   17.10
15 $                 85 0.665691 $                   56.58
16 $                 85 0.647874 $                   55.07
17 $                 85 0.630535 $                   53.60
18 $                 85 0.613659 $                   52.16
19 $                 85 0.597235 $                   50.76
20 $                 85 0.581251 $                   49.41
20 $           1,000 0.581251 $                 581.25
Bond Price $             1,049.55

Correct Option: THIRD